OpenStudy (anonymous):

Solve the system of equation using the addition method. 3x-4y=3 2x=5y-2

OpenStudy (anonymous):

First, line up up each variable and symbol, then choose a variable you want to eliminate. Multiply each term in one equation by something which will make the variables cancel out when you add the equations together. Then, add the equations together and you're left with one variable you can solve for.

OpenStudy (anonymous):

that didn't really help

OpenStudy (campbell_st):

rewrite your equations 3x - 4y = 3 (1) 2x - 5y = -2 (2) multiply equation(1) by 2 and add to equation (2) times - 3 to eliminate x 6x - 8y = 6 -6x + 15y = 6 -------------- 7y = 12 solve for y...